# Exponentiations of ultrafilters

### v3

@@ 2   Exponentiations of ultrafilters
First of all, since $E_{1},E_{2}$ are not associative, we cannot approach the existence of idempotents problem via the usual straightforward application of Ellis’ Theorem (see [5, Theorem 2.5.] for details). Moreover, since $1$ is the only principal $E_{1}-$idempotent, and the only $E_{2}$-idempotent as well, from now on we will use “idempotent” to actually mean “nonprincipal idempotent”.
In the following we want to study properties of generalizations of the operation $(p,q)\rightarrow 2^{p}\odot q$, first introduced by Di Nasso and Ragosta in [2]. We will use the following notations: for $n\in\mathbb{N},p,q\in\beta\mathbb{N}$ we let:
$p\oleft_{n}q:=n^{p}\odot q;p\oright_{n}q:=p\odot n^{q}.$
The following result shows that the existence of idempotents for $E_{2}$ is related to that of ⊣⃝$\oleft_{n}⊣$-idempotents, whilst the existence of idempotents for $E_{1}$ is related to that of ⊢⃝$\oright_{n}⊢$-idempotents.
Proposition 2.15.
If $n\in\mathbb{N}_{>1}$, then $(1)\Rightarrow(2)\Rightarrow(3)$, where
  1. (i)
    there exists $p\in\beta\mathbb{N}$ such that $p\oplus n^{p}=p$;
  2. (ii)
    there exists $p\in\beta\mathbb{N}$ such that $p\oright_{n}p=p$;
  3. (iii)
    there exists $p\in\beta\mathbb{N}$ such that $E_{1}(p,p)=p$.
Analogously, $(4)\Rightarrow(5)\Rightarrow(6)$, where
  1. (iv)
    there exists $p\in\beta\mathbb{N}$ such that $n^{p}\oplus p=p$;
  2. (v)
    there exists $p\in\beta\mathbb{N}$ such that $p\oleft_{n}p=p$;
  3. (vi)
    there exists $p\in\beta\mathbb{N}$ such that $E_{2}(p,p)=p$.
Proof.
All the implications follows by taking $q$ as given by the hypothesis, by letting $p=n^{q}$ and by applying Proposition 2.4. For this reason, we prove only that $(v)$ entails $(vi)$, as the other implications are very similar.
$(v)\Rightarrow(vi)$ Let $q\in\beta\mathbb{N}$ be such that $q\oleft_{n}q=q$ and let $p=n^{q}$. Then $E_{2}(p,p)=E_{2}\left(n^{q},n^{q}\right)$. As $n^{q}=E_{2}(q,n)$, by Proposition 2.4.(6) we have that
$E_{2}\left(n^{q},n^{q}\right)=E_{2}\left(n^{q},E_{2}\left(q,n\right)\right)=E_{2}\left(n^{q}\odot q,n\right)=E_{2}\left(q,n\right)=p.\qed$
Moreover, in the search of $E_{1}$- and $E_{2}$-idempotent ultrafilters, we make the following observation:
@@
We just have to apply Proposition 2.4.(5)-(6).
In fact, if $p=p\odot E_{1}(q,p)$ then
$E_{1}(q,p)=E_{1}(q,p)=E_{1}(q,p\odot E_{1}(q,p))=E_{1}\left(E_{1}(q,p),E_{1}(q,p)\right)$
@@
All that has been said so far might point to the actual existence of exponential idempotents. Actually, as we are going to show, this is not the case for $E_{1}$, and at least not trivially the case for $E_{2}$.
As for $E_{1}$, the non existence of idempotents follows from an even stronger fact, that we now prove.
@@
By [6, Lemma 23], there is a $4$-coloring $\mathbb{N}=A_{1}\cup\dots\cup A_{4}$ of $\mathbb{N}$ such that there is no monochromatic pair $b,a^{b}$ with $\log_{2}a\leq b$ and $a,b\geq 1$.
@@
Now take $A_{i}\in\{A_{1},\dots,A_{4}\}$ such that $A_{i}\in q$. In particular, $B=\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A_{i}\}\in q\}\neq\emptyset$, so we can pick $n\in B$. Let $I_{n}:=\{m\in\mathbb{N}\mid n^{m}\in A_{i}\}$. As $q$ is non principal, $I_{n}$ must contain some $m>\log_{2}(n)$. But $m,n^{m}\in A_{i}$ with $\log_{2}(n)\leq m$, which is absurd. ∎
Putting together Theorem 2.17 and Propositions 2.15, 2.16, we get the following:
@@
For all $n>1\in\mathbb{N}$ and $p,q\in\beta\mathbb{N}\setminus\{0,1\}$ we have that $p\oplus n^{p}\neq p$ and $p\odot E_{1}(q,p)\neq p$.
Notice that, in particular, also $p\oright_{n}{p}=p$ has no solution in $\beta\mathbb{N}$ if $n>1$.
@@
By contrast, let $p$ be such an ultrafilter. Let $S:\mathbb{N}\rightarrow\mathbb{N}$ be the function that maps $n=p_{1}^{a_{1}}\cdot\dots\cdot p_{k}^{a_{k}}$ to $\sum_{i\leq k}a_{i}$, where $n=p_{1}^{a_{1}}\cdot\dots\cdot p_{k}^{a_{k}}$ is the prime decomposition of $n$. Observe that $S(E_{2}(p_{1},p_{2}))=p_{1}\odot S(p_{2})$ and $S(p_{1}\odot p_{2})=S(p_{1})\oplus S(p_{2})$ for all $p_{1},p_{2}\in\beta\mathbb{N}$. In particular, $S(p)=S(E_{2}(p,p))=p\odot S(p)$ and $S(p)=S(p\odot p)=S(p)\oplus S(p)$, therefore $S(p)$ is additively idempotent and $p\odot S(p)=S(p)\oplus S(p)$. If $S(p)\in\mathbb{N}$, then by additive idempotency we would get that $S(p)=0$, from which $p=1$, which is absurd as $p\in\beta\mathbb{N}\setminus\mathbb{N}$. So $S(p)$ is additively idempotent and nonprincipal, so $n\mathbb{N}\in S(p)$ for all $n\in\mathbb{N}$. Therefore we can apply [5, Theorem 13.14] to deduce that it must be $p\odot S(p)\neq S(p)\oplus S(p)$, reaching the desired contrast. ∎
@@ 3   Open questions
Let $p\in E\left(K(\beta\mathbb{N},\odot)\right)$ be such that each member of $p$ is also additive central. Is it true or not that each $A\in p$ contains a configuration of the form $\{m,n,m+n,m2^{n}\}$?
@@
Are there $p,q\in\beta\mathbb{N}\setminus\mathbb{N}$ that solve any of the following equations
@@
  7. 7.
    $E_{2}(q,p)\odot q=q$.
@@
Question 3.3.
Does it exist $p$ that is $E_{2}$-idempotent? Does it exist $p$ that is ⊣⃝$\oleft_{n}⊣$-idempotent?
@@
Question 3.4.
What are the algebraical properties of $EVdW$, besides Theorem 2.12?

### v4

@@ 2   Exponentiations of ultrafilters
First of all, since $E_{1},E_{2}$ are not associative, we cannot approach the existence of idempotents problem via the usual method of applying Ellis’ Theorem (see [5, Theorem 2.5] for details). Moreover, since $1$ is the only principal $E_{1}-$idempotent, and the only $E_{2}$-idempotent as well, from now on we will use “idempotent” to actually mean “nonprincipal idempotent”. Our starting observation is the following:
@@
We just have to apply Proposition 2.10.(5)-(6). In fact, if $p=p\odot E_{1}(q,p)$ then
$E_{1}(q,p)=E_{1}(q,p\odot E_{1}(q,p))=E_{1}\left(E_{1}(q,p),E_{1}(q,p)\right)$
@@
In particular, the existence of $n\in\mathbb{N}_{>1}$ and $p\in\beta\mathbb{N}\setminus\mathbb{N}$ with $p\odot n^{p}=p$ would force $n^{p}$ to be an $E_{1}$-idempotent; similarly, if $n^{p}\odot p=p$ then $n^{p}$ would be $E_{2}$-idempotent.
However, idempotents cannot exist for $E_{1}$, and cannot exist for $E_{2}$ under certain additional assumptions, as we are now going to show. As for $E_{1}$, the non existence of idempotents follows from an even stronger fact, that we now prove.
@@
By [7, Lemma 23], there is a $4$-coloring $\mathbb{N}=A_{1}\cup\ldots\cup A_{4}$ of $\mathbb{N}$ such that there is no monochromatic pair $b,a^{b}$ with $\log_{2}a\leq b$ and $a,b\geq 1$.
@@
Now take $A_{i}\in\{A_{1},\ldots,A_{4}\}$ such that $A_{i}\in q$. In particular, $B=\{n\in\mathbb{N}\mid\{m\in\mathbb{N}\mid n^{m}\in A_{i}\}\in q\}\neq\emptyset$, so we can pick $n\in B$. Let $I_{n}:=\{m\in\mathbb{N}\mid n^{m}\in A_{i}\}$. As $q$ is non principal, $I_{n}$ must contain some $m>\log_{2}(n)$. But $m,n^{m}\in A_{i}$ with $\log_{2}(n)\leq m$, which is a contradiction. ∎
Putting together Theorem 2.26 and Proposition 2.25, we get the following:
@@
For all $n>1\in\mathbb{N}$ and $p,q\in\beta\mathbb{N}\setminus\{0,1\}$ we have that $p\odot n^{p}\neq p$ and $p\odot E_{1}(q,p)\neq p$.
@@
By contrast, let $p$ be such an ultrafilter. Let $\Omega:\mathbb{N}\rightarrow\mathbb{N}$ be the function that maps $n$ to the sum of the exponents in its prime factorization. In particular, $\hat{\Omega}(p)=\hat{\Omega}(E_{2}(p,p))=p\odot\hat{\Omega}(p)$ and $\hat{\Omega}(p)=\hat{\Omega}(p\odot p)=\hat{\Omega}(p)\oplus\hat{\Omega}(p)$, therefore $\hat{\Omega}(p)$ is additively idempotent and $p\odot\hat{\Omega}(p)=\hat{\Omega}(p)\oplus\hat{\Omega}(p)$. If $\hat{\Omega}(p)\in\mathbb{N}$, then by additive idempotency we would get that $\hat{\Omega}(p)=0$, from which $p=1$, which is a contradiction as $p\in\beta\mathbb{N}\setminus\mathbb{N}$. So $\hat{\Omega}(p)$ is additively idempotent and nonprincipal, hence $n\mathbb{N}\in\hat{\Omega}(p)$ for all $n\in\mathbb{N}$. Thus we can apply [5, Theorem 13.14] to deduce that $p\odot\hat{\Omega}(p)\neq\hat{\Omega}(p)\oplus\hat{\Omega}(p)$, reaching the desired contradiction. ∎
@@ 3   Open questions
Let $p\in E\left(K(\beta\mathbb{N},\odot)\right)$ be such that each member of $p$ is also additively central. Is it true or not that each $A\in p$ contains a configuration of the form $\{m,n,m+n,m2^{n}\}$?
@@
Prove that there exist $p,q,r\in\beta\mathbb{N}\setminus\mathbb{N}$ that solve any of the following equations
@@
  8. 8.
    $E_{2}(q,p)\odot q=q$;
  9. 9.
    $E_{2}(p,r)\odot E_{2}(q,r)=E_{2}(p\oplus q,r)$.
For one ultrafilter, the following question, for which we conjecture a negative answer, seems relevant:
Question 3.3.
Does it exist $p\in\beta\mathbb{N}\setminus\mathbb{N}$ such that $E_{1}(p,p)=E_{2}(p,p)$?
@@
Question 3.4.
Does it exist $p$ that is $E_{2}$-idempotent?
@@
Question 3.5.
What are the algebraical properties of $\mathrm{ESW}$?